### v1

# Diffusion in driven periodic systems: from weak to strong damping regimes

Abstract
We consider diffusion of an inertial Brownian particle moving in a symmetric spatially periodic potential and driven by an unbiased time periodic force of a moderate frequency. The study is focused on the dependence of its diffusion coefficient $D$ on the damping $\gamma$ ranging from the weak to strong one. We reveal the mechanisms of counterintutive effects in which diffusion can increase despite the grow of friction and exhibits both minima and maxima. Their origin is explained by analysis of the deterministic dynamics perturbed by thermal fluctuations. In the weak damping regime the maximum of diffusion is induced by a non-monotonic change of population of the running orbits when the friction increases. In the strong dissipation regime the effect is caused by bifurcation of periodic orbits oscillating symmetrically around the potential maximum and their displacement towards the potential minimum when the damping coefficient increases. It induces the change of the effective barrier of the periodic potential and its vanishing at the maximum of $D$.
I Introduction
The celebrated Einstein relation for the diffusion coefficient $D$ of the force-free Brownian particle is given by the following relationship einstein ; entropy23
$D=\frac{k_{B}T}{\gamma},$ (1)
where $k_{B}$ is the Boltzmann constant, $T$ is temperature of the medium (thermostat) in which the particle moves and $\gamma$ is the friction (damping) coefficient which characterizes the interaction strength of the Brownian particle with the environment. The same relation is also fulfilled for a Brownian particle subjected to a constant force $F={const}$. From this equation it follows that $D$ depends inversely on the friction coefficient $\gamma$. In other words, $D$ decreases when $\gamma$ increases. Studies on e.g. surface diffusion reveals that in some regimes the more general relation is satisfied, namely,
$D\sim\frac{1}{\gamma^{\sigma}}.$ (2)
The exponent $\sigma$ is not universal but depends on a system, its model and the corresponding assumptions. Different values have been reported like $\sigma=1/2$ sigma05 , $\sigma=1/3$ sigma03 , $\sigma\in(0.64,0.76)$ sigma064 , $\sigma\in(0,1/3)$ sigma033 . Non-standard behavior of $D$ on the damping coefficient $\gamma$ is also possible. A good example is Brownian diffusion of skyrmions within magnetic materials skyrmion1 . For vanishing topological charge, $D$ monotonically decreases with growing $\gamma$. However, if the topological charge is non-zero, the diffusion coefficient increases until it reaches its maximum and next it decreases when the damping constant grows skyrmion2 . In the paper, we study the issue of non-monotonic dependence of diffusion on the friction coefficient for the Brownian particle moving in a spatially periodic potential and driven by a time-periodic force, i.e., the model which is relevant to a variety of different context risken , including pendulums gitterman , Josephson junctions kautz1996 , superionic conductors fulde , dipoles rotating in external fields coffey , Frenkel-Kontorova models frenkel , charge density waves gruner and cold atoms dwelling in optical lattices renzoni ; denisov , to name only a few. We mention that a similar system, however with an additional constant force, has been studied in Ref. march-prl .
The layout of the present work is as follows. In Sec. II, we describe the model of the Brownian particle moving in a spatially periodic potential driven by an unbiased time-periodic force. It is defined in terms of the dimensionless Langevin equation. Its full dimensional version is presented in Appendix A. In Sec. III, we analyze the diffusion coefficient as a function of dissipation, i.e. how $D$ depends on the friction coefficient $\gamma$. We show that there are several regimes in which (i) $D$ monotonically decreases with respect to $\gamma$; (ii) $D$ is non-monotonic and exhibits maximum as a function of $\gamma$; (iii) there are many extrema of $D$ as $\gamma$ grows. In Sec. IV, we implement the vibrational mechanics approximation march-prl which is valid for high frequency of the ac-driving force and examine this approximation for the analysis of diffusion coefficient as a function of damping. Section V provides a summary and conclusions.
II Model of Brownian dynamics
We study diffusion properties of the Brownian particle moving in a spatially periodic structure and subjected to a time-periodic force. The corresponding Langevin equation assumes the dimensionless form marchPRE23
$\ddot{x}+\gamma\dot{x}=-\sin{x}+a\sin(\omega t+\phi_{0})+\sqrt{2\gamma Q}\,\xi(t).$ (3)
Note that in this scaling the dimensionless mass $m=1$. The parameter $\gamma$ is the friction coefficient and $Q\propto k_{B}T$ is the dimensionless temperature of the system. The coupling of the particle with thermostat is modeled by the $\delta$-correlated Gaussian white noise $\xi(t)$ of zero mean and unit intensity, i.e. $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(s)\rangle=\delta(t-s)$. The starting dimensional equation is presented in Appendix A, where the corresponding scaling and dimensionless parameters are defined.
The complexity of the underlying dynamics with the nonlinear force $f(x)=-\sin{x}$, the time-periodic force $g(t)=a\sin(\omega t+\phi_{0})$ of the amplitude $a$, frequency $\omega$, initial phase $\phi_{0}$ and thermal fluctuations of the intensity $2\gamma Q$ do not allow for a analytical approach which presently is clearly beyond the known mathematical methods. Therefore numerical simulations have been employed spiechowicz2015cpc .
In the deterministic case, when $Q=0$, the system may be non-ergodic spiechowicz2016scirep and therefore sensitive to the specific choice of the initial conditions: position $x(0)$, velocity $v(0)$ of the particle and the initial phase $\phi_{0}$ of the ac-driving. Therefore all quantities characterizing the system should be averaged over $\{x(0),v(0),\phi_{0}\}$ with uniform distributions to get rid of this dependence. However, for any non-zero temperature $Q>0$ the system is ergodic and the initial conditions do not affect its properties in the long time stationary regime. Therefore the mean value of observables can be accurately estimated by averaging any realization over a sufficiently long time interval.
III Diffusion coefficient
Despite its apparent simplicity, the system (3) is extremely versatile, exhibits an impressive diversity of behavior and can successfully describe a number of important and various physical situations. We will be interested only in one aspect of these properties, namely, diffusion and its dependence on dissipation characterized by the friction coefficient $\gamma$. Because the parameters space $\{\gamma,a,\omega,Q\}$ is four dimensional we consider only the selected parameter regime in which the diffusion process in the long time limit is normal and described by a time-independent diffusion coefficient $D$. Hence it can be defined as spiechowicz2016scirep
$D=\lim_{t\to\infty}\frac{1}{2t}\,\langle\left[x(t)-\langle x(t)\rangle\right]^{2}\rangle.$ (4)
III.1 Diffusion versus friction: monotonic decrease
The exact expression for $D$ is known in several cases simpler than the current one. For the force-free Brownian particle or under the influence of a constant force $F$, i.e. when Eq.(3) assumes the form
$\ddot{x}+\gamma\dot{x}=F+\sqrt{2\gamma Q}\,\xi(t),$ (5)
the diffusion coefficient it given by the Einstein relation,
$D=\frac{Q}{\gamma}.$ (6)
For the Brownian particle moving in the periodic structure, i.e. when the Langevin equation reads
$\ddot{x}+\gamma\dot{x}=-\sin{x}+\sqrt{2\gamma Q}\,\xi(t),$ (7)
the corresponding diffusion coefficient is derived in two limiting damping regimes in which it can be expressed as
$D=\frac{Q}{\gamma}\,B(Q),$ (8)
where $B(Q)$ depends only on the dimensionless temperature $Q$.
In the underdamped regime ($\gamma\ll 1$), the factor $B(Q)$ reads pavli ; entropy23
$B(Q)=\frac{\sqrt{\pi/Q}}{2I_{0}(1/Q)}\int_{1}^{\infty}\frac{\mbox{e}^{-y/Q}\,dy}{\sqrt{y+1}\,\mathbf{E}(\sqrt{2/(y+1)})},$ (9)
where $I_{0}(x)$ is the modified Bessel function of the first kind and $\mathbf{E}(k)$ is the complete elliptic integral of the second kind specialF .
On the other hand, in the opposite case of overdamped dynamics, i.e. when the inertial term in Eq. (7) vanishes, it assumes the following form lifson ; entropy23
$B(Q)=\frac{1}{I_{0}^{2}(1/Q)}.$ (10)
In the above three cases, the diffusion coefficient $D$ depends inversely on the friction coefficient $\gamma$, i.e., $D\sim 1/\gamma$. In other words, if the damping is stronger then the diffusion coefficient is smaller.
Figure 1: Dimensionless diffusion coefficient $D$ as a function of the amplitude $a$ of the ac-driving for its fixed frequency ${\omega=1.59}$ and two values of the friction coefficient $\gamma$. Temperature reads $Q=0.05$.
Figure 2: Dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$ for the ac-driving amplitude $a=6$ and its frequency $\omega=1.59$. The temperature is $Q=0.5$.
When the ac-force is included as in Eq. (3), the analytical form of $D$ is not known and only numerical results are available. In Ref. march-chaos it was shown that for a tailored parameter regime the diffusion coefficient $D$ displays the damped quasi-periodic dependence on the amplitude $a$ of the time-periodic force $g(t)=a\sin{(\omega t+\phi_{0})}$, as an example see Fig. 1. The origin of this phenomenon was explained in Ref. marchPRE23 and is rooted in the complex deterministic structure of the nonlinear dynamics governed by a variety of stable and unstable orbits.
We remind that the running and localized states (periodic orbits) can coexist in this system marchPRE23 . The particle can oscillate in the periodic potential wells and next proceed either forward or backward within one or many temporal periods, or it can permanently move in one direction. There are several types of contribution to the spread of trajectories and in consequence to the diffusion coefficient $D$. The first, which is the largest one, comes from the distance between the running trajectories moving into the opposite directions. The second, also relatively large, is the spread between the running and periodic orbits. The next follows from contributions driven by thermal fluctuation in each of the running and localized states. Clearly, even at high temperature the latter are remarkably smaller than the two previous ones. Therefore it is intuitively expected that the diffusion coefficient $D$ will change considerably especially when the population of running and locked states is significantly modified.
The dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$ has not been studied in Ref. marchPRE23 . In Fig. 1 we show the influence of $\gamma$ on the behavior of $D$ for low temperature regime $Q=0.05$. It is seen that $D$ decreases as $\gamma$ increases, which portrays the expected standard behavior. So, it is similar to the previous cases presented in Eqs. (6) and (8).
In Fig. 2 we show the characteristics $D$ versus $\gamma$ for a set of fixed parameters. For the driving amplitude $a=6$ and temperature $Q=0.5$ (the first local maximum of $D$ in Fig. 1) one can distinguish three distinct intervals of functional dependence: (i) for damping $\gamma\in(0.01,0.1)$ the diffusion coefficient diminishes like $1/\gamma^{1.27}$; (ii) for $\gamma\in(0.1,8)$, it vanishes like $1/\gamma^{2}$; (iii) for larger $\gamma$ it again decreases like $1/\gamma$. The cases (i) and (ii) show different behavior than this in Eqs. (6) and (8). For the driving amplitude $a=10$ (the first local minimum of $D$ in Fig. 1), one can observe a gentle slope in the vicinity of $\gamma\sim 0.5$ and $1/\gamma$ is detected for $\gamma\gg 1$ (not depicted). The above exemplified monotonic dependence of $D$ on $\gamma$ occurs in extended domain of parameter space, in particular for large values of $\gamma$ and high temperature $Q$. However, this picture is changed radically when temperature $Q$ is significantly lowered.
III.2 Diffusion versus friction: non-monotonic dependence
The case illustrated in Fig. 2 is in fact only a quantitative modification of the monotonic decay $D\sim 1/\gamma$. Much more interesting is the non-monotonic dependence $D(\gamma)$. Such a situation is a qualitatively new phenomenon: the non-monotonicity implies the occurrence of local extrema, i.e. maxima and/or minima. An example of this behavior is visualized in Fig. 3. For the driving amplitude $a=14$ corresponding to the second local maximum of $D$ in Fig. 1 and lower temperature $Q=0.05$, the diffusion coefficient decreases for small values of $\gamma$. Next, it attains the local minimum and starts to increase to the local maximum at $\gamma\approx 0.06$. If $\gamma$ continues to grow, the diffusion coefficient monotonically diminishes to the next local minimum and increases up to the next maximum at $\gamma\approx 3.345$. Further increase of $\gamma$ results in a monotonic decline of $D$ to zero. The main question is why $D$ starts to grow when the friction increases and whether this behavior is exception or can occur in a wider domain of the parameters space. Let us note that the local maxima can be detected for both weak ($\gamma\approx 0.06$) as well as strong damping (${\gamma\approx 3.345}$). However, the mechanism behind this effect is radically different in both cases.
At this place it is worth mentioning that in the system there are localized states $V=0$ (the particle oscillates in one or several potential wells and the motion is bounded in the coordinate space) and running states $V=\pm k\omega,k=1,2,3,...$, where $\omega$ is the dimensionless frequency of the periodic force and the period averaged velocity is defined as
[…]
Figure 3: Non-monotonic dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$ for the ac-driving amplitude $a=14$ (the second maximum in Fig. 1). The frequency of the ac-driving reads $\omega=1.59$ and temperature is $Q=0.05$.
Figure 4: Stationary probability P of the Brownian particle to reside in various states $V=0,V=\pm k\omega\}$ characterizing the deterministic counterpart of the studied setup. The black line corresponds to the sum of probabilities for all running states. The amplitude is $a=14$ and the frequency reads $\omega=1.59$. Temperature is $Q=0.05$.
In the insets of Fig. 3 the probability distribution $P(V)$ for the period-averaged velocity $V$ of the Brownian particle is depicted for the detected local maxima of diffusion coefficient $D$. For the weak damping, in the vicinity of $\gamma=0.06$, the probability distribution has three remarkable maxima, one corresponding to the locked state $V=0$ and two symmetric running states of the averaged velocities $V=\pm\omega$. At non-zero temperature, there are random transitions induced by thermal fluctuations between these states. For one trajectory, the particle can follow the running solution, next for some time interval it resides in the localized state and later it again travels along the running state either in the same or the opposite direction. In this way, for a given trajectory various deterministic solutions are visited in a random sequence. In consequence the deterministic structure of locked and running states impacts the spread of trajectories.
In Fig. 4 we consider the probability P to reside in various states $\{V=0,\pm k\omega\}$ characterizing the deterministic counterpart of the studied setup. It allows us to answer the question about the mechanism standing behind the observed pronounced maximum of the diffusion coefficient $D$ as a function of the friction $\gamma$ for weak damping regime $\gamma\approx 0.06$. The initial decay of this characteristics is clearly related to a decrease of the running states $V=\pm k\omega$ and simultaneous grow of the locked ones $V=0$. Next, there is one-to-one correspondence between the non-monotonic behavior of $D$ and the probability P to reside in the running states. It means that the local growth of the number of running trajectories is responsible for the increase of $D$ despite the fact that the damping $\gamma$ is getting bigger as well.
Figure 5: Probability density $p(x)$ for the position of the particle in the long time regime averaged over the initial conditions $x_{0}$, $v_{0}$ and $\phi_{0}$. The system is deterministic $Q=0$. The potential $U(x)=-\cos{x}$ is depicted schematically for the convenience of the reader. Other parameters are: $a=14$, $\omega=1.59$ and the damping coefficient: $(a)\,\gamma=3.15;$ $(b)\,\gamma=3.21;$ $(c)\,\gamma=3.3;$ $(d)\,\gamma=3.345;$ $(e)\,\gamma=3.36;$ $(f)\,\gamma=3.41$.
As it is well evidenced in Fig. 4 for strong damping regime $\gamma>1$ the running states cease to exist and the period-averaged velocity distribution $P(V)$ exhibits only one maximum corresponding to the locked orbits, c.f. Fig. 3. Yet we can still detect a non-monotonic behavior of the diffusion coefficient $D$ as a function of the friction coefficient $\gamma$ with a well pronounced maximum observed for $\gamma_{M}\approx 3.345$. Therefore the previously outlined mechanism of this effect for low to moderate $\gamma$ is no longer valid in this regime. We shall now explain the origin of this phenomenon which is clearly different than the former one.
We start our consideration assuming the value of $\gamma=3.15<\gamma_{M}$: if $\gamma$ increases then $D$ increases. To explain this effect we first look at the probability density $p(x)$ for the position of the particle in the long time regime. In Fig. 5 we present this quantity for the deterministic trajectories $Q=0$ averaged with respect to the initial position $x_{0}$, velocity $v_{0}$ and phase $\phi_{0}$ distributed uniformly over the intervals $[0,2\pi]$, $[-2,2]$ and $[0,2\pi]$, respectively. We note that the averaging over the phase $\phi_{0}$ in the long time regime is equivalent to the averaging over the period of the external driving. For the convenience of the reader in the same plot we additionally illustrate the periodic potential $U(x)=-\cos{x}$. The reader can observe the characteristic arc-sine form of $p(x)$ levy ; arcsine which here is a consequence of averaging over the phase $\phi_{0}$.
It turns out that in the presented regime corresponding to $\gamma=3.15$ (panel (a) in Fig. 5) the deterministic particle symmetrically oscillates around the maximum $x=x_{M}=\pi$ of the potential and covers the indicated distance $\Delta x$ between its minimal and maximal position. The quantity $\Delta x/2$ can be interpreted as the amplitude of oscillations and the boundary values of this interval correspond to the most probable values of the particle position. Oscillations around the maximum of the potential (the inverted pendulum) are not astonishing taking into account the large amplitude $a=14$ of the time periodic force. Note that the total ”rocking” potential $V(x,t)=-\cos(x)-ax\cos(\omega t+\phi_{0})$ has local extrema (and the barriers) only when $a|\cos(\omega t+\phi_{0})|<1$. It means that for the amplitude $a=14$ in majority of time intervals ($95.5\%$ of full time) there are no barriers. Only in very, very short time itervals ($4.5\%$ of full time) the barriers emerge and the potential is indeed the washboard one. Two extreme forms $V_{1}=-\cos{x}-14x$ and $V_{2}=-\cos{x}+14x$ looks like straight lines, the barriers disappear for some time intervals and the particle can ”dance” around any position.
If now $\gamma$ grows to the value $\gamma_{1}=3.2$, the amplitude of oscillations decreases and the spatial interval $\Delta x$ around the maximum $x_{M}$ of the potential $U(x)$ decreases. It means that in this window the effective potential barrier the particle needs to overcome to diffuse due to thermal agitation is reduced as well. It explains why the diffusion coefficient $D$ increases despite the grow of damping $\gamma$.
Next, at $\gamma_{1}=3.2$ the ”phase transition” arises: two periodic orbits emerge, oscillate around the potential maximum corresponding to the indicated intervals $\Delta x_{1}$ and $\Delta x_{2}$ and they move in the opposite directions when $\gamma$ increases, see panel (c) in Fig. 5. These oscillations are not symmetric around $x_{M}$ but their centers are shifted towards the potential minimum. It means that one of the turning points $x_{e}$ of the orbit (where the particle velocity changes sign) is closer to the potential maximum and the barrier height is further reduced, see the two inner vertical blue and red lines in panel (c) of Fig. 5.
For $\gamma_{M}=3.345$ one of the turning points of both blue and red orbits is exactly at the potential maximum $x_{M}$ and consequently the barrier height is zero, cf. panel (d) in Fig. 5. When the particle is approaching $x_{M}$ its velocity decreases to zero and for a relatively long time it resides in the neighborhood of $x_{M}$. In consequence, thermal fluctuations can freely spread the particle and diffusion is maximal.
For the friction coefficient lying in the interval $\gamma\in(3.345,3.365)$ one observes two periodic orbits related to non-symmetrical oscillations around the potential minima for which the effective barrier height is greater than for $\gamma_{M}$ and therefore the diffusion coefficient starts to decrease as it is expected, see panel (e) in Fig. 5. At $\gamma_{2}=3.365$ the reversed bifurcation is detected: two previously observed locked trajectories now symmetrically oscillate around two neighbouring potential minima and can be regarded as the same by a shift of the blue orbit to the left, see panel (f) in Fig. 5.
For $\gamma>\gamma_{2}$ we observe a decrease in the spatial interval $\Delta x$ of oscillations around the minimum. In other words, oscillations are focused closer and closer to the minimum of the potential as $\gamma$ grows. As a result, the effective potential barrier height increases and the diffusion coefficient decreases. To sum it up, the sequence of panels in Fig. 5 reveals the mechanism which is responsible for the non-monotonic behavior of diffusion coefficient $D$ in the strong damping regime.
IV Vibrational mechanics approach
The diffusional behavior for strong damping can be satisfactory explained in terms of the effective potential barrier for the particle motion in a spatially periodic potential. To this aim, we use vibrational mechanics methods and replace the starting Eq. (3) by the following Langevin equation boroEPL ; march-prl
$\ddot{\tilde{x}}+\gamma\dot{\tilde{x}}=-J_{0}(\psi)\sin{\tilde{x}}+\sqrt{2\gamma Q}\,\xi(t),$ (12)
where $\tilde{x}=\tilde{x}(t)$ is a slow part of the dynamics (the fast one is represented by free oscillations) and $J_{0}(\psi)$ is the zero order Bessel function specialF of the argument
$\psi=\frac{a}{\omega\sqrt{\omega^{2}+\gamma^{2}}}.$ (13)
We note that the starting total time-dependent potential $V(x,t)=-\cos(x)-ax\sin(\omega t+\phi_{0})$ is replaced by the time-independent renormalized potential $V_{eff}(x)=-J_{0}(\psi)\cos(x)$. In this approach, the particle diffuses in a renormalized potential of the effective barrier $2J_{0}(\psi)$. This approximation is postulated to be correct for the high frequency limit $\omega\gg 1$. In this section, we want to verify to what extend this approximation reconstructs the exact dynamics. In Fig. 6 we compare results obtained from the exact dynamics (3) and its approximate counterpart (12).
For the case of weak dissipation, we can apply the modified equation (9) which reads pavli ; entropy23
$D=\frac{Q}{\gamma}\,\frac{\sqrt{\pi d_{0}/Q}}{2I_{0}(d_{0}/Q)}\int_{1}^{\infty}\frac{\mbox{e}^{-d_{0}y/Q}\,dy}{\sqrt{y+1}\,\mathbf{E}(\sqrt{2/(y+1)})},$ (14)
where $d_{0}=J_{0}(\psi)$. In panel (b) of Fig. 6 we compare the dependence $D(\gamma)$ from this relation with the exact results of simulations. The discrepancy is very big and (14) cannot reconstruct the local maximum for $a=14$ which is at $\gamma=0.06$. Therefore this approximation totally fails.
However, for strong damping, the approximate expression for the diffusion coefficient can be obtained from the modified equation (10) which takes the following form lifson ; entropy23
$D=\frac{Q}{\gamma I_{0}^{2}(J_{0}(\psi)/Q)}.$ (15)
Figure 6: Diffusion $D$ versus the friction coefficient $\gamma$ depicted for the 1st (panel (a)) and the 2nd (panel (b)) maximum and minimum of $D$ taken from Fig. 1 together with a comparison of the numerical results and the approximate formula given by Eq. (14) and (15). The ac-driving frequency reads $\omega=1.59$ while temperature is $Q=0.05$.
In this case, the approximation (15) correctly reproduces the exact results for the friction coefficient $\gamma>0.3$ which could be treated as the regime of strong damping. Moreover, it turns out that for larger values of the amplitude $a$ new maxima emerge, e.g. for $a=18$ there are two maxima at $\gamma_{M}=1.05$ and $\gamma_{M}=1.4$ but for $a=35$ four maxima exist at $\gamma_{M}=0.95,2,3.75,9.5$ (not shown). The envelope of these points behaves roughly like $D(\gamma_{M})\sim 1/\gamma_{M}$.
V Summary
We showed that the dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$ characterizing the dissipation mechanism can render unexpected features. Typically, diffusion $D$ is a decreasing function of the damping $\gamma$. In contrast, we demonstrated that for an archetypal model of nonequilibrium system consisting of a Brownian particle moving in a symmetric and spatially periodic potential subjected to an unbiased time-periodic driving the diffusion coefficient $D$ can grow when the damping $\gamma$ increases. Moreover, for sufficiently large amplitude of the ac-driving the characteristics $D(\gamma)$ can exhibit several local maxima and minima in the strong damping regime.
We revealed the mechanisms standing behind this counterintuitive phenomenon. In the weak damping limit the maximum of diffusion is induced by a non-monotonic change of population of the running states when the dissipation increases. In contrast, in the strong friction regime it is related to vanishing of the effective barrier of the potential in which the Brownian particle dwells. The latter effect is caused by displacement of the location (center) and amplitude of the locked orbits induced by the damping change. Properties in the strong dissipation regime are well reproduced by the vibrational mechanics methods proposed in Refs. boroEPL ; march-prl .
Acknowledgments
This work was supported by the Grants NCN 2022/45/B/ST3/02619 (J.S.) and NASU No. 0122U002145/2022-2023 (V.A). I. G. M. acknowledges University of Silesia for its hospitality since the beginning of the war, 24 February 2022.
Appendix A Nondimensionalization of the Langevin equation
The starting dimensional Langevin equation has the form entropy23 ; marchPRE23 ; spiechowicz2016njp
$M\ddot{x}+\Gamma\dot{x}=-U^{\prime}(x)+F(t)+\sqrt{2\Gamma k_{B}T}\,\xi(t),$ (16)
where $M$ is the mass of the Brownian particle, $U(x)$ is the spatially periodic potential $U(x)=U(x+L)$ of period $L$ and $F(t)$ is an unbiased and symmetric time-periodic force. Moreover, the dot and the prime denotes differentiation with respect to time $t$ and the particle coordinate $x$, respectively. The parameter $\Gamma$ is the friction (damping) coefficient. The potential $U(x)$ has the simplest form
$U(x)=-\Delta U\cos{\left(\frac{2\pi}{L}x\right)},$ (17)
where $2\Delta U$ is the barrier height. The external ac-driving force of amplitude $A$ and angular frequency $\Omega$ is taken in the form
$F(t)=A\sin{(\Omega t)}.$ (18)
Thermal equilibrium fluctuations describing the influence of thermostat of temperature $T$ are normally assumed as $\delta$-correlated Gaussian white noise of zero-mean value,
$\langle\xi(t)\rangle=0,\quad\langle\xi(t)\xi(s)\rangle=\delta(t-s),$ (19)
[…]
We define the following dimensionless coordinate and time
$\hat{x}=2\pi\frac{x}{L},\quad\hat{t}=\frac{t}{\tau_{0}},\quad\tau_{0}=\frac{L}{2\pi}\sqrt{\frac{M}{\Delta U}}.$ (20)
The characteristic length $L$ is the period of the potential $U(x)$ and the characteristic time $\tau_{0}$ is related to the period of small oscillations inside the potential well of $U(x)$.
Under the above scaling, Eq. (16) is converted to the form
$\ddot{\hat{x}}+\gamma\dot{\hat{x}}=-\sin{\hat{x}}+a\sin(\omega\hat{t})+\sqrt{2\gamma Q}\,\hat{\xi}(\hat{t}),$ (21)
where the dimensionless parameters read
$\gamma=\frac{\Gamma}{M}\tau_{0},\quad a=\frac{1}{2\pi}\frac{L}{\Delta U}A,\quad\omega=\tau_{0}\Omega,\quad Q=\frac{k_{B}T}{\Delta U}.$
The dimensionless potential $V(x)$ of the period $L=2\pi$ is $V(\hat{x})=U((L/2\pi)\hat{x})/\Delta U=-\cos{\hat{x}}$ and the corresponding potential force reads $f(x)=-\hat{U}^{\prime}(\hat{x})=-\sin\hat{x}$. The rescaled thermal noise is $\hat{\xi}(\hat{t})=(L/2\pi\Delta U)\xi(t)=(L/2\pi\Delta U)\xi(\tau_{0}\hat{t})$ and has the same statistical properties as $\xi(t)$; i.e., $\langle\hat{\xi}(\hat{t})\rangle=0$ and $\langle\hat{\xi}(\hat{t})\hat{\xi}(\hat{s})\rangle=\delta(\hat{t}-\hat{s})$. The dimensionless noise intensity $Q$ is the ratio of thermal energy and half of the activation energy the particle needs to overcome the nonrescaled potential barrier. In order to simplify the notation we omit the hat-notation in Eq. (3).